OpenStudy (scarlettfarra2000):

Please Help! I medal solve the following equations show all work for full credit

OpenStudy (tigga281):

Whats the question?

OpenStudy (scarlettfarra2000):

\[7v ^{2}+1=29\]

OpenStudy (tigga281):

Give me two mins

OpenStudy (scarlettfarra2000):

\[9x ^{2}-1=63\]

OpenStudy (tigga281):

Let's solve your equation step-by-step. 7v2+1=29 Step 1: Subtract 1 from both sides. 7v2+1−1=29−1 7v2=28 Step 2: Divide both sides by 7. 7v2 7 = 28 7 v2=4 Step 3: Take square root. v=±√4 v=2 or v=−2 Answer: v=2 or v=−2

OpenStudy (tigga281):

Let's solve your equation step-by-step. 9x2−1=63 Step 1: Add 1 to both sides. 9x2−1+1=63+1 9x2=64 Step 2: Divide both sides by 9. 9x2 9 = 64 9 x2= 64 9 Step 3: Take square root. x=±√ 64 9 x=2.666667,−2.666667 Answer: x=2.666667,−2.666667

OpenStudy (tigga281):

Anymore? I have to go soon

OpenStudy (anonymous):

9x^2-1=63 +1 +1 9x^2=64 = 8/3 or -8/3

OpenStudy (anonymous):

I believe that is the answer but let me check real quick to make sure.

OpenStudy (scarlettfarra2000):

I no more I only had two thank you for your help

OpenStudy (tigga281):

No problem anytime just tag me

OpenStudy (tigga281):

If it's in fraction form then it's the same answer as mine, just turn the decimal into a fraction